### v4

@@ 4.2 Main Results on Deepfake Generation
Table 7: Results of representative face swapping methods on FF++. Notations: ➊ CelebA-HQ, ➋ FFHQ, ➌ VGGFace, ➍ VGGFace2, ➎ VoxCeleb2.
@@
FaceShifter li2019faceshifter ➊➋➌ 97.38 2.06 2.96 -
SimSwap chen2020simswap 92.83 - 1.53 -
FaceInpainter li2021faceinpainter ➊➋➌ 97.63 - 2.21 -
HifiFace wang2021hififace 98.48 - 2.63 -
RAFSwap xu2022region 96.70 2.92 2.53 -
Xu et alxu2022high 90.05 2.79 2.46 -
DiffSwap zhao2023diffswap 98.54 5.35 2.45 2.16
FlowFace zeng2023flowface ➊➋➍ 99.26 - 2.66 -
FlowFace++ zhang2023flowface++ ➊➋➍ 99.51 - 2.20 -
StyleIPSB jiang2023styleipsb 95.05 2.23 3.58 -
StyleSwap xu2022styleswap ➌➎ 97.05 5.28 1.56 2.72
WSC-Swap ren2023reinforced ➊➋➌ 99.88 5.01 1.51 -
Table 8: Results of representative face reenactment methods on VoxCeleb for the self-reenactment. Notations: ➊ VoxCeleb, ➋ VoxCeleb2, ➌ ETH-Xgaze, ➍ Gaze360 kellnhofer2019gaze360 , ➎ MPIIGaze zhang2015appearance , ➏ TalkingHead-1KH. In addition, we use gray to represent data that is partially uncertain.
@@
HyperReenact bounareli2023hyperreenact 0.710 - 0.230 27.10 -
DG hsu2022dual 0.831 - - 22.10 0.761
AVFR-GAN agarwal2023audio - 32.20 - 08.48 0.824
Free-HeadGAN doukas2023free ➊➌➍➎ 0.810 22.16 0.100 35.40 -
HiDe-NeRF li2023one ➊➋➐ 0.931 21.90 0.084 - 0.862
Table 9: Results of representative face reenactment methods on VoxCeleb for the cross-identity reenactment. Notations: ➊ VoxCeleb, ➋ VoxCeleb2, ➌ ETH-Xgaze, ➍ Gaze360 kellnhofer2019gaze360 , ➎ MPIIGaze zhang2015appearance , ➏ TalkingHead-1KH.
@@
HyperReenact bounareli2023hyperreenact 0.680 - - - -
AVFR-GAN agarwal2023audio - - - 09.05 -
Free-HeadGAN doukas2023free ➊➌➍➎ 0.789 - - 53.90 13.1
HiDe-NeRF li2023one ➊➋➏ 0.786 0.012 0.971 57.00 -
Table 10: Results of representative face reenactment methods on VoxCeleb2 for quality assessment. Notations: ➊ VoxCeleb, ➋ VoxCeleb2, ➌ LRW, ➍ CelebV-HQ, ➎ TalkingHead-1KH.
@@
PC-AVS zhou2021pose ➋➌ - - 6.880 - 0.886
GC-AVT liang2022expressive - - 2.757 - 0.739
Wang et alwang2023memory - 28.92 1.830 - 0.830
PECHead gao2023high ➋➍➎ 1.590 - 23.05 -
DG hsu2022dual 0.721 - - 51.79 0.540
HiDe-NeRF li2023one ➊➎ 0.787 - - 61.00 -
$\bullet$ Results on Talking Face Generation. Table 11 displays the performance results of various talking face generation approaches on the MEAD wang2020mead dataset since 2023. AMIGO zhai2023talking achieves promising results among all methods, which utilizes seq2seq to generate facial landmarks for emotion tagging, matching mouth movements with emotional features. Additionally, it employs a landmark-to-image translation network to create facial images with fine textures.
$\bullet$ Results on Facial Attribute Editing.  Table 12 evaluates the quality level of generated images using FID, and Table 13 assesses facial reconstruction capabilities using PSNR, LPIPS, and SSIM. Due to different training and testing datasets, quantitative fair comparisons are not possible that just serve as performance demonstrations.
Table 11: Evaluation results of the models involved in talking face generation on the MEAD dataset. Notations: ➊ MEAD, ➋ LRW, ➌ VoxCeleb2, ➍ HDTF.
@@
Xu et alxu2023high 0.83 2.36 - 3.500 15.91 30.09/0.850
EMMN tan2023emmn ➊➋ - - 2.780/2.870 3.570 - 29.38/0.660
AMIGO zhai2023talking ➊➌ - 2.44 2.140/2.440 - 19.59 30.29/0.820
SLIGO sheng2023stochastic 0.88 1.83 - 3.690 - 00000-/0.790
Gan et algan2023efficient ➊➌ - - 2.250/2.470 - 19.69 21.75/0.680
DreamTalk ma2023dreamtalk ➊➌➍ - - 2.910/1.930 3.780 - 00000-/0.860
SPACE gururani2023space ➊➌ - - - 3.610 11.68 0-
TalkCLIP ma2023talkclip - - 3.601/2.415 3.773 - 00000-/0.829
Table 12: FID evaluation of different methods. Notations: ➊ FFHQ, ➋ CelebA-HQ, ➌ MM CelebA-HQ, ➍ CelebA, ➎ CelebAText-HQ, ➏ Self-build.
@@
FENeRF sun2022fenerf GANs+NeRF 12.10
FENeRF sun2022fenerf GANs+NeRF 28.20
AnyFace sun2022anyface GANs 56.75
AnyFace sun2022anyface GANs 50.56
TextFace hou2022textface GANs 22.81
TG-3DFace yu2023towards GANs 52.21
TG-3DFace yu2023towards GANs 39.02
HifaFace gao2021high GANs ➊➋ ➊➋ 04.04
GuidedStyle hou2022guidedstyle GANs 41.79
4.3 Main Results on Forgery Detection
The mainstream evaluation metrics for forgery detection technology are ACC and AUC. Table 14 presents the ACC and AUC metrics for some detection models trained on FF++ rossler2019faceforensics++ and tested on FF++ (HQ) and FF++ (LQ). LipForensics haliassos2021lips exhibits robust performance on the strongly compressed FF++ (LQ), while Guo et alguo2023controllable perform best on FF++ (HQ). Table 15 shows the cross-dataset evaluation using DFDC dolhansky2020deepfake , Celeb-DF li2020celeb , Celeb-DFv2 li2020celeb , and DeeperForensics-1.0 jiang2020deeperforensics as validation sets. AVoiD-DF yang2023avoid and Zhao et alzhao2021learning demonstrate excellent generalization ability, but there is still significant room for improvement in these datasets. However, overall, there is room for improvement in the evaluation performance of forgery detection models on the DFDC, which has a larger sample size and more complex forgery methods.
Table 13: Results of facial reconstruction capabilities in facial attribute editing work. Notations: ➊ FFHQ, ➋ VoxCeleb, ➌ VoxCeleb2, ➍ CelebA-HQ.
Methods Type Train Test PSNR↑ LPIPS↓ SSIM↑
Preechakul et alpreechakul2022diffusion Difussion - 0.0110 0.991
Kim et al. kim2023diffusion Difussion - 0.0450 0.922
FDNeRF zhang2022fdnerf GANs+NeRF - 0.1420 0.821
IA-FaceS huang2023ia GANs ➊➍ 22.34 0.2240 0.642
IA-FaceS huang2023ia GANs ➊➍ 22.43 0.0384 0.659
Table 14: Results of the self-dataset performance on FF++. Notations: HQ (Mild compression), LQ (Heavy compression).
@@ 4.3 Main Results on Forgery Detection
F3-Net qian2020thinking FF++ 93.02 95.80 98.95 99.30
Masi et almasi2020two FF++ 86.34 - 96.43 -
Zhao et alzhao2021multi FF++ 88.69 90.40 97.60 99.29
FDFL li2021frequency FF++ 89.00 92.40 96.69 99.30
LipForensics haliassos2021lips FF++ 94.20 98.10 98.80 99.70
RECCE cao2022end FF++ 91.03 95.02 97.06 99.32
Guo et alguo2023controllable FF++ 92.76 96.85 99.24 99.75
MRL yang2023masked FF++ 91.81 96.18 93.82 98.27
Table 15: Results of cross-dataset performance evaluation on four datasets DFDC, Celeb-DF (CDF), Celeb-DFv2 (CDFv2), and DeeperForensics-1.0 (DFo). Evaluation indicator is AUC. Notations: ➊ FF++, ➋ FF++(Real), ➌ FF++(HQ), ➍ FF++(LQ), ➎ Self-build, ➏ SR-DF wang2022m2tr , ➐ DFDCp, ➑ FakeAVCeleb, ➒ DefakeAVMiT.
@@
Face X-ray li2020face ➊➎ 80.92 80.58 - -
Face X-ray li2020face 71.15 74.76 - -
Masi et almasi2020two - 73.41 - -
Zhao et alzhao2021multi 67.44 - - -
Zhao et alzhao2021learning 67.52 98.30 90.03 99.41
Zheng et alzheng2021exploring 74.00 - 86.90 98.80
LipForensics haliassos2021lips 73.50 - 82.40 97.60
M2TR wang2022m2tr - 82.10 - -
M2TR wang2022m2tr - 68.20 - -
RECCE cao2022end 69.06 68.71 - -
SBIs shiohara2022detecting - - 90.79 -
SBIs shiohara2022detecting 72.42 - 93.18 -
RealForensics haliassos2022leveraging 75.90 - 86.90 99.30
Guo et alguo2023controllable 81.65 84.97 - -
Yin et alyin2023dynamic 73.08 71.36 - -
MRL yang2023masked 71.53 83.58 - -
AVoiD-DF yang2023avoid 80.60 - - -
AVoiD-DF yang2023avoid 90.30 - - -
5 Future Prospects
This section provides a short discussion of future research directions that can facilitate and envision the development of deepfake generation and detection.
@@ 5 Future Prospects
$\bullet$ Talking Face Generation. Current methods strive to enhance the realism of generated conversational videos. However, they lack fine-grained control over the emotional nuances of the conversation, where the matching of emotional intonation to audio and semantic content is not precise enough, and the control over emotional intensity is too coarse. In addition, the realistic correlation between head pose and facial expression movement seems insufficiently considered. Lastly, for text or audio with intense emotions, noticeable artifacts in head movement still occur to a significant extent.
@@
6 Conclusion
@@
Acknowledgement. This work is supported by the National Natural Science Foundation of China (No. 62371189).

### v5

@@ 4.2. Benchmark Protocol
Table 7. Results of representative face swapping methods on FF++. Notations: ➊ CelebA-HQ, ➋ FFHQ, ➌ VGGFace, ➍ VGGFace2, ➎ VoxCeleb2.
@@
FaceShifter (li2019faceshifter, ) ➊➋➌ 97.38 2.06 2.96 -
SimSwap (chen2020simswap, ) 92.83 - 1.53 -
HifiFace (wang2021hififace, ) 98.48 - 2.63 -
RAFSwap (xu2022region, ) 96.70 2.92 2.53 -
Xu et al(xu2022high, ) 90.05 2.79 2.46 -
DiffSwap (zhao2023diffswap, ) 98.54 5.35 2.45 2.16
FlowFace (zeng2023flowface, ) ➊➋➍ 99.26 - 2.66 -
StyleIPSB (jiang2023styleipsb, ) 95.05 2.23 3.58 -
StyleSwap (xu2022styleswap, ) ➌➎ 97.05 5.28 1.56 2.72
WSC-Swap (ren2023reinforced, ) ➊➋➌ 99.88 5.01 1.51 -
DiffFace (kim2025diffface, ) - 2.71 2.35 -
CanonSwap (luo2025canonswap, ) 99.78 - 1.59 -
Table 8. Results of representative face reenactment methods on VoxCeleb for the self-reenactment. Notations: ➊ VoxCeleb, ➋ VoxCeleb2, ➌ ETH-Xgaze, ➍ Gaze360 (kellnhofer2019gaze360, ), ➎ MPIIGaze (zhang2015appearance, ), ➏ TalkingHead-1KH. In addition, we use gray to represent data that is partially uncertain.
@@
HyperReenact (bounareli2023hyperreenact, ) 0.710 - 0.230 27.10 -
DG (hsu2022dual, ) 0.831 - - 22.10 0.761
AVFR-GAN (agarwal2023audio, ) - 32.20 - 08.48 0.824
Free-HeadGAN (doukas2023free, ) ➊➌➍➎ 0.810 22.16 0.100 35.40 -
HiDe-NeRF (li2023one, ) ➊➋➐ 0.931 21.90 0.084 - 0.862
DiffusionAct (bounareli2025diffusionact, ) ➊➋ 0.690 19.70 0.240 - 0.830
Table 9. Results of representative face reenactment methods on VoxCeleb for the cross-identity reenactment. Notations: ➊ VoxCeleb, ➋ VoxCeleb2, ➌ ETH-Xgaze, ➍ Gaze360 (kellnhofer2019gaze360, ), ➎ MPIIGaze (zhang2015appearance, ), ➏ TalkingHead-1KH.
@@
HyperReenact (bounareli2023hyperreenact, ) 0.680 - - - -
AVFR-GAN (agarwal2023audio, ) - - - 09.05 -
Free-HeadGAN (doukas2023free, ) ➊➌➍➎ 0.789 - - 53.90 13.1
HiDe-NeRF (li2023one, ) ➊➋➏ 0.786 0.012 0.971 57.00 -
DiffusionAct (bounareli2025diffusionact, ) ➊➋ 0.600 - - - -
Table 10. Results of representative face reenactment methods on VoxCeleb2 for quality assessment. Notations: ➊ VoxCeleb, ➋ VoxCeleb2, ➌ LRW, ➍ CelebV-HQ, ➎ TalkingHead-1KH.
@@
PC-AVS (zhou2021pose, ) ➋➌ - - 6.880 - 0.886
GC-AVT (liang2022expressive, ) - - 2.757 - 0.739
Wang et al(wang2023memory, ) - 28.92 1.830 - 0.830
PECHead (gao2023high, ) ➋➍➎ 1.590 - 23.05 -
DG (hsu2022dual, ) 0.721 - - 51.79 0.540
HiDe-NeRF (li2023one, ) ➊➎ 0.787 - - 61.00 -
$\bullet$ Forgery Detection. This task requires evaluation under both self-dataset and cross-dataset settings, with AUC and ACC as the commonly used metrics. In this survey, self-dataset performance is evaluated on the FF++ dataset, with results separately reported for FF++ (HQ) and FF++ (LQ) in terms of AUC and ACC, as shown in  Tab. 14. For cross-dataset evaluation, four representative datasets—DFDC, Celeb-DF, Celeb-DFv2, and DeeperForensics-1.0—are selected, and the corresponding AUC results are reported in  Tab. 15.
4.3. Main Results on Deepfake Generation
$\bullet$ Results on Face Swapping. Tab. 8 displays the performance evaluation results of some representative models on the Face Swapping task using the FF++ (rossler2019faceforensics++, ) dataset. WSC-Swap (ren2023reinforced, ) captures external facial attribute information and internal identity features through two independent encoders, enabling strong identity preservation and stable facial pose retention. However, it shows sub-optimal performance on facial expression error metrics. In addition, the method has been reported to suffer from certain attribute loss during facial attribute transfer.
$\bullet$ Results on Face Reenactment. Tab. 8 and Tab. 10 show the performance evaluation results on the VoxCeleb (nagrani2017voxceleb, ) dataset for self-reenactment and cross-subject reenactment, respectively. AVFR-GAN (agarwal2023audio, ) achieves better performance by using the multimodal modeling.  Tab. 10 presents quality assessment results on the VoxCeleb2 dataset. HiDe-NeRF (li2023one, ) represents 3D scenes using canonical appearance fields and implicit deformation fields. It achieves accurate facial attribute modeling by explicitly decoupling facial pose and expression attributes using a deformation module.
Table 11. Evaluation results of the models involved in talking face generation on the MEAD dataset. Notations: ➊ MEAD, ➋ LRW, ➌ VoxCeleb2, ➍ HDTF.
@@ 4.3. Main Results on Deepfake Generation
Xu et al(xu2023high, ) 0.83 2.36 - 3.500 15.91 30.09/0.850
EMMN (tan2023emmn, ) ➊➋ - - 2.780/2.870 3.570 - 29.38/0.660
AMIGO (zhai2023talking, ) ➊➌ - 2.44 2.140/2.440 - 19.59 30.29/0.820
SLIGO (sheng2023stochastic, ) 0.88 1.83 - 3.690 - 00000-/0.790
Gan et al(gan2023efficient, ) ➊➌ - - 2.250/2.470 - 19.69 21.75/0.680
SPACE (gururani2023space, ) ➊➌ - - - 3.610 11.68 0-
TalkCLIP (ma2025talkclip, ) - - 3.601/2.415 3.773 - 00000-/0.829
MTHM (tang2025one, ) - - - - 22.38 21.38/0.660
Table 12. Results of facial reconstruction capabilities in representative facial attribute editing work. Notations: ➊ FFHQ, ➋ VoxCeleb, ➌ VoxCeleb2, ➍ CelebA-HQ.
Methods Type Train Test PSNR↑ LPIPS↓ SSIM↑
Konpat et al(preechakul2022diffusion, ) Difussion - 0.0110 0.991
Kim et al. (kim2023diffusion, ) Difussion - 0.0450 0.922
FDNeRF (zhang2022fdnerf, ) GANs+NeRF - 0.1420 0.821
HairNeRF (chang2023hairnerf, ) GANs+NeRF 31.84 0.1060 0.827
IA-FaceS (huang2023ia, ) GANs ➊➍ 22.34 0.2240 0.642
IA-FaceS (huang2023ia, ) GANs ➊➍ 22.43 0.0384 0.659
r-FaceS (deng2024r, ) GANs ➊➍ 30.69 0.0220 -
$\bullet$ Results on Talking Face Generation. Tab. 12 displays the performance results of various talking face generation approaches on the MEAD (wang2020mead, ) dataset since 2023. AMIGO (zhai2023talking, ) achieves promising results, which utilizes seq2seq to generate facial landmarks for emotion tagging, matching mouth movements with emotional features. Additionally, it employs a landmark-to-image translation network to create facial images with fine textures.
$\bullet$ Results on Facial Attribute Editing.  Tab. 12 evaluates the quality level of generated images using FID, and Tab. 14 assesses facial reconstruction capabilities using PSNR, LPIPS, and SSIM. Due to different training and testing datasets, quantitative fair comparisons are not possible that just serve as performance demonstrations.
Table 13. FID evaluation of different methods. Notations: ➊ FFHQ, ➋ CelebA-HQ, ➌ $M$ CelebA-HQ, ➍ CelebA, ➎ CelebAText-HQ, ➏ Self-build.
@@
FENeRF (sun2022fenerf, ) GANs+NeRF 12.10
FENeRF (sun2022fenerf, ) GANs+NeRF 28.20
AnyFace (sun2022anyface, ) GANs 56.75
AnyFace (sun2022anyface, ) GANs 50.56
TextFace (hou2022textface, ) GANs 22.81
TG-3DFace (yu2023towards, ) GANs 52.21
TG-3DFace (yu2023towards, ) GANs 39.02
HifaFace (gao2021high, ) GANs ➊➋ ➊➋ 04.04
GuidedStyle (hou2022guidedstyle, ) GANs 41.79
NeRFFaceEditing (jiang2025towards, ) GANs 06.00
Table 14. Results of the self-dataset performance on FF++. HQ (Mild compression), LQ (Heavy compression).
@@
F3-Net (qian2020thinking, ) FF++ 93.02 95.80 98.95 99.30
Masi et al(masi2020two, ) FF++ 86.34 - 96.43 -
Zhao et al(zhao2021multi, ) FF++ 88.69 90.40 97.60 99.29
FDFL (li2021frequency, ) FF++ 89.00 92.40 96.69 99.30
LipForensics (haliassos2021lips, ) FF++ 94.20 98.10 98.80 99.70
RECCE (cao2022end, ) FF++ 91.03 95.02 97.06 99.32
Guo et al(guo2023controllable, ) FF++ 92.76 96.85 99.24 99.75
MRL (yang2023masked, ) FF++ 91.81 96.18 93.82 98.27
Table 15. Results of cross-dataset performance evaluation on four datasets DFDC, Celeb-DF (CDF), Celeb-DFv2 (CDFv2), and DeeperForensics-1.0 (DFo). Evaluation indicator is AUC. Notations: ➊ FF++, ➋ FF++(Real), ➌ FF++(HQ), ➍ FF++(LQ), ➎ Self-build, ➏ SR-DF (wang2022m2tr, ), ➐ DFDCp, ➑ FakeAVCeleb, ➒ DefakeAVMiT.
@@
Face X-ray (li2020face, ) ➊➎ 80.92 80.58 - -
Zhao et al(zhao2021learning, ) 67.52 98.30 90.03 99.41
Zheng et al(zheng2021exploring, ) 74.00 - 86.90 98.80
LipForensics (haliassos2021lips, ) 73.50 - 82.40 97.60
M2TR (wang2022m2tr, ) - 82.10 - -
M2TR (wang2022m2tr, ) - 68.20 - -
RECCE (cao2022end, ) 69.06 68.71 - -
SBIs (shiohara2022detecting, ) - - 90.79 -
SBIs (shiohara2022detecting, ) 72.42 - 93.18 -
RealForensics (haliassos2022leveraging, ) 75.90 - 86.90 99.30
Guo et al(guo2023controllable, ) 81.65 84.97 - -
Yin et al(yin2023dynamic, ) 73.08 71.36 - -
MRL (yang2023masked, ) 71.53 83.58 - -
AVoiD-DF (yang2023avoid, ) 80.60 - - -
WMamba (peng2025wmamba, ) 82.97 96.29 - -
4.4. Main Results on Forgery Detection
Tab. 14 presents the ACC and AUC metrics for some detection models trained on FF++ (rossler2019faceforensics++, ) and tested on FF++ (HQ) and FF++ (LQ). LipForensics (haliassos2021lips, ) exhibits robust performance on the strongly compressed FF++ (LQ), while Guo et al(guo2023controllable, ) perform best on FF++ (HQ). Tab. 15 shows the cross-dataset evaluation. AVoiD-DF (yang2023avoid, ) and Zhao et al(zhao2021learning, ) demonstrate excellent generalization ability, but there is still significant room for improvement in these datasets. However, overall, there is room for improvement in the evaluation performance of forgery detection models on the DFDC.
5. Future Prospects
@@
$\bullet$ Talking Face Generation. Current methods aim to improve the realism of generated conversational videos, yet they still lack fine-grained control over emotional dynamics. The alignment between emotional intonation and audio–semantic content remains imprecise, and the modulation of emotional intensity is overly coarse. Moreover, the realistic coupling between head pose and facial expression is often under-modeled. Finally, for text or audio conveying strong emotions, existing approaches still produce noticeable artifacts in head motion.
@@
$\bullet$ Discussion. Future development of deepfake generation and detection technologies will move toward more robust, adaptive and multimodal systems. For generation, integrating visual, auditory and textual cues can strengthen semantic consistency and identity preservation, while larger and higher-quality datasets will support better cross-domain generalization. Reinforcement learning and feedback-driven optimization may further enhance visual fidelity and temporal coherence. Detection research is expected to benefit from multimodal evidence aggregation and finer temporal modeling, leveraging cues such as audio–video synchronization, speech semantics and physiological signals. Self-supervised and weakly supervised learning will also help reduce dependence on extensive annotations as new manipulation types continue to emerge. Deepfake generation and detection will expand into applications including digital humans, telepresence, assistive media creation and privacy-preserving data synthesis. As adoption grows, ethical risks related to identity misuse, privacy violations and unauthorized content generation become more pressing. Incorporating transparency mechanisms such as watermarking and provenance metadata, along with strong misuse detection and clear governance guidelines, will be essential for ensuring the safe and responsible deployment of deepfake technologies.
6. Conclusion